A 25.5-g object moving to the right at 20.5 cm/s overtakes and collides elastically with a 12.0-g object moving in the same direction at 15.0 cm/s. Find the velocity of each object after the collision. (Take the positive direction to be to the right. Indicate the direction with the sign of your answer.)

25.5-g object     cm/s
12.0-g object     cm/s
Answer #1

Suppose after collision 12g is moving with v1 and 25.5g object with v2 in same direction,

collision is elastic, so

velocity of approach = velocity of separation

20.5 - 15 = v1 - v2

v1 = 5.5 + v2 .........(i)

Using momentum conservation for this collision,

25.5 x 20.5 + 12 x 15 = 25.5v2 + 12v1

12v1 + 25.5v2 = 702.75

putting value of v1 from eqaution(i)

12(5.5 + v2) + 25.5v2 = 702.75

37.5v2 = 636.75

v2 = 16.98 cm/s ....................velocity of 25.5 g object.

v1 = 5.5 + v2 = 22.48 cm/s .........velocity of 12 g object.